Suppose that 132.0 J of heat energy is required to completely vaporize 3.1 g of a pure liquid. If the molar mass of the liquid is 40.17 g/mol, what is the molar enthalpy of vaporization of the liquid?
1
Expert's answer
2019-09-16T03:55:18-0400

n= m/M = 3.1 / 40.17 =0.07717 mol

132 / 0.07717 = 1710.51 J/mol